# ReMA: Learning to meta-think for LLMs with multi-agent reinforcement learning

### v2

@@ 3.2 Training MAMRP: A Multi-agent RL Method
where $R_{h}$ and $R_{l}$ are policies’ individual reward functions, including $R$ and regularization according to tasks and models, e.g., different format rewards. In this work, we conduct experiments to enhance MAMRP described in Equation 7 using the update method outlined in Equations 10 and 11. For the detailed design of our experiments, we’ll discuss different designs of reward functions in Sec 3.3. With the above objective functions, we apply multi-agent RL algorithms to enhance collaboration between the two models. Additionally, to stabilize the training process, we conduct prompt filtering as described in Cui et al. (2025): during each round of training, we filter out the questions that our models solve with a success rate less than $\varepsilon_{\min}$ or greater than $\varepsilon_{\max}$. Similar to Kirchner et al. (2024b), we train the models iteratively with RL algorithms as shown in the Algorithm 1. We illustrate the MAMRP inference procedure and the proposed training method in Fig. 2. We also provide an analysis of different loss functions in Appendix F.
3.3 Reward Function Designs
As described in Section 3.2, we update both high-level and low-level agents by assigning rewards based on the low-level policy output. Below, we outline several potential reward designs:
The combination of different rewards enhances the overall performance of LLMs. Detailed reward designs are presented in Appendix E.
@@ 4 Experiments
To evaluate the effectiveness and efficiency of ReMA, we conduct experiments on challenging benchmarks for two types of tasks: mathematical reasoning and LLM-as-a-Judge, and we train and test different LLMs.
@@
4.1 Benchmarks
For mathematical reasoning experiments, we train models on 7.5k training samples in MATH (Hendrycks et al., 2021) and use MATH500 (Lightman et al., 2023) as the in-distribution test dataset. Additionally, we test the optimized models on out-of-distribution datasets: GSM8K (Cobbe et al., 2021), AIME24111https://huggingface.co/datasets/AI-MO/aimo-validation-aime, AMC23222https://huggingface.co/datasets/AI-MO/aimo-validation-amc, GaoKao2023En (Zhang et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad Bench (He et al., 2024).
For LLM-as-a-Judge benchmarks, we train models on RewardBench (Lambert et al., 2024). Specifically, we convert the original data into a pair-ranking format with two possible orders, creating 5.97k items. We split them into a training set of 5k items and a test set of 970 items, denoted as RewardBench970. The models are also tested on JudgeBench (Tan et al., 2024) to assess out-of-distribution performance. We’ve tested two different split strategies, and we refer to Appendix B for detailed comparisons and results.
4.2 Baselines, Models, and Training Details
We compare performance across the following methods:
  • VRP (CoT): Prompting the LLM to solve the problem step-by-step (VRP in Sec 3.1).
[…]
  • ReMA (ours): Using RL to optimize the LLMs’ under MAMRP as defined in Eq 7.
And for each method, we show the greedy decoding performance in Section 4.3. We train and test Llama-3-8B-Instruct, Llama-3.1-8B-Instruct (Dubey et al., 2024), and Qwen2.5-7B-Instruct (Team, 2024) on mathematical reasoning benchmarks. For LLM-as-a-judge benchmarks, we train and test Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct. We use instruct-tuned LLMs so as to prompt them to perform VRP, MRP and MAMRP during training. Unless specified, we use two copies of the same model for high- and low-level agent in ReMA. We use the base reward setting in Sec E by default. For the underline RL algorithm, we use REINFORCE++ (Hu, 2025a). We refer to Appendix I for detailed training hyper-parameters and training settings.
4.3 Main Results
Table 1 compares the greedy decoding performance of ReMA against various RL baselines across mathematical reasoning benchmarks (Table 1(a)) and LLM-as-a-Judge benchmarks (Table 1(b)). Results across different LLMs indicate that, on average, ReMA outperforms all baselines, achieving a maximum average improvement of 6.68% on mathematical benchmarks and 8.49% on LLM-as-a-Judge benchmarks.
Notably, ReMA achieves the highest performance on most benchmarks, particularly on out-of-distribution datasets, with a maximum improvement of 20% on AMC23 for Llama3-8B-Instruct, 13.33% on AIME24 for Qwen2.5-7B-Instruct, 14.23% on RewardBench970 for Llama3.1-8B-Instruct. However, we observe that the accuracy gains from RL training on instruction-tuned models are smaller than from base models (Sec 4.4.1). Beyond the higher initial performance of instruction-tuned models, we hypothesize that this is due to their relatively fixed output distribution, which limits their adaptability and exploration in reinforcement learning. Additionally, the relatively small model size and dataset scale may further constrain the improvements achieved through reinforcement learning.
4.4 Ablations
@@ 4.4.1 Meta-thoughts boost low-level generalization
Figure 3: An RL experiment evaluates the performance of Qwen2.5-Math-7B on the MATH500, GSM8K, and AIME24 datasets after training on Level 3-5 MATH questions using different methods. RL from SFT achieves superior performance but struggles to generalize to more challenging problems. In contrast, RL from Base and RL under Meta-thinking demonstrate the ability to solve previously unseen, harder problems, with the latter further enhancing performance.
@@
Here we provide a minimal but motivating example of how ReMA gives better learning dynamics. We use Qwen2.5-Math-7B (Yang et al., 2024) as the starting base model, MATH (level 3-5, about 5.5K number of instances) as the training dataset, and we compare three reinforcement learning training schemes, in particular:
  1. 1.
    RL from Base: train the base model directly on MATH with binary outcome reward;
  2. 2.
    RL from SFT: First, collect a small amount of CoT data (from GPT-4o’s answers) to fine-tune the base model (SFT); then train the tuned model on MATH with binary outcome reward;
  3. 3.
    RL under Meta-thinking: First, collect a small amount of high-level meta-thinking data (from GPT-4o’s plans) to fine-tune Qwen2.5-Math-7B-Instruct (Yang et al., 2024) to increase its meta-thinking abilities and make sure it act as high-level planner only; then train Qwen2.5-Math-7B as the low-level agent in collaboration with the high-level instruction with binary outcome reward.
The model is tested on a testing set of MATH500, GSM8K, and AIME24. The results are shown in Figure 3. First, supervised fine-tuning (SFT) largely boosts performance as it has the best accuracy on the in-distribution test set MATH500 and the easier one GSM8K at the beginning of RL training. However, it doesn’t show advances on the harder benchmark AIME24.
More importantly, while RL from SFT significantly boosts performance, it struggles to generalize to harder problems. In contrast, RL from Base shows limited improvements, especially on more challenging tasks. Notably, RL under Meta-thinking achieves the best learning dynamics, effectively leveraging high-level guidance to improve the performance of the lower-level solving agent, particularly on harder problems like AIME2024. We refer to Appendix J.1 for a detailed case study.
4.4.2 Reward functions shape cross-agent behaviors
We also investigate the impact of different reward function designs on ReMA’s behavior. In addition to the base reward setting described in Sec 3.3, we evaluate a consistency-based reward function using Qwen2.5-7B-Instruct. This reward function is designed to encourage the high-level agent to generate more detailed guidance. Indeed, we observe that the high-level agent trained in this manner produces more detailed solution steps compared to the one trained with the basic correctness format reward. However, we also find that this approach often leads to jailbreak behavior, where the high-level agent tends to include the final answer within its output, compromising the intended hierarchical reasoning process.
This suggests that (1) meta-thinking can not only emerge and be reinforced in the high-level agent but also in the low-level agent. During reinforcement learning (RL) training, the two agents develop a novel problem-solving pattern characterized by a role reversal. (2) Consistency-based rewards promote a more self-corrective approach at the lower level, potentially disrupting the intended separation of roles between planning and execution. For a detailed case study, refer to Appendix J.2.
4.5 Diverse meta-thinking characteristics of LLMs
Figure 4: Average problem difficulty of three predefined metacognitive actions during training. Left: After approximately 20 training steps, the 1B LM’s outputs collapse to the simplest action, EMPTY, resulting in no data points for “REWRITE” and “DECOMPOSE” in the plot. Right: In contrast, the 8B LM learns to utilize more complex metacognitive actions for solving difficult problems.
To further analyze the meta-thinking behaviors learned by ReMA, we conduct a simple interpretability experiment through training with predefined JSON-format action sets inspired by Yue et al. . Specifically, we modify the high-level agent’s decision-making process by first selecting from three predefined meta-thinking actions: DECOMPOSE (breaking the problem into subproblems), REWRITE (reformulating the problem to simplify it), and EMPTY (solving the problem directly). After selecting an action, the agent then generates the corresponding text for that action within a single LM call, formatting the output in JSON.
We compare two experimental setups using Llama-3.1-8B-Instruct and Llama-3.2-1B-Instruct. These configurations allow us to study models of different scales: one setup consists of two 1B models, while the other consists of two 8B models. This setup enables us to investigate the impact of RL on the final training results of the high-level agent. To ensure the model maintains valid output formatting, we employ VLLM guided JSON decoding (Kwon et al., 2023; Dong et al., 2024) during rollout. We use the base reward setting, i.e., the averaged accuracy of the low-level agent’s solutions with format constraints.
First, we observe that the outputs of smaller language models (LMs) tend to converge to relatively simpler and shorter sentences compared to those in our main experiments. This may be due to the limited capacity of smaller LMs to generate valid JSON-formatted responses while simultaneously exploring diverse reasoning strategies. As shown in Fig. 4, Llama-3.2-1B-Instruct struggles to capture the complex dynamics underlying the task. Instead, it quickly converges to outputting only the EMPTY action to minimize penalties associated with formatting errors. In contrast, Llama-3.1-8B-Instruct successfully learns to employ different metacognitive strategies based on the difficulty level of the problem. A detailed case study is provided in Appendix J.3.
5 Related work
Drawing from the bitter lesson (Sutton, 2019), two methods that appear to scale effectively are searching and learning, aligning with current trends in large language models (Xu et al., 2025). At present, researchers are leveraging these methods to maximize the capabilities of individual transformers, while other efforts are exploring architectures that involve multiple interacting entities. In this paper, we examine this divergence within the context of LLM reasoning, a capability that allows large language models to solve problems through logical reasoning, step-by-step analysis, and inference (Wang et al., 2024a).
5.1 Single LLM Reasoning
Main research works in reasoning involving a single LLM utilize search-based and post-training methods. The fundamental elements of searching methods are text generation and evaluation. Generation schemes include In-Context Learning (Brown et al., 2020), Beam Search (Graves, 2012), and various tree-based searching (Snell et al., 2024); Evaluation approaches often use outcome accuracy, self-consistency (Wang et al., 2022), or process reward signal (Lightman et al., 2023) as the criteria to select high-quality responses from the generated texts. Post-training method is another research line in opposition to pre-training. Popular training pipelines often involve specific data construction followed by Supervised Fine-tuning (Qin et al., 2024; Ouyang et al., 2022; Hui et al., 2024; Liu et al., 2024), or reinforcement learning to interactively explore learning patterns (Wang et al., 2024a; Zhang et al., 2024a; DeepSeek-AI et al., 2025; Xu et al., 2025).
5.2 Multiple LLM Reasoning
Integrating multiple entities can potentially surpass the intelligence of the individual model (Chen et al., 2023). With the rapid emergence of large language models showing a varying level of abilities, some studies have explored facilitating discussions among multiple off-the-shelf LLMs (Zhang et al., 2025a; Chen et al., 2024a; Wang et al., 2023; Du et al., 2023; Zhuge et al., 2023; Tang et al., 2023; Hao et al., 2025; Akata et al., 2023; Hong et al., 2023; Zhang et al., 2024b), taking the form of free discussion (Du et al., 2023; Liang et al., 2023) or structured role assignments (Hong et al., 2023; Zhang et al., 2024b). Some have applied routing mechanisms to assign tasks to the most suitable expert models (Hu et al., 2024b; Stripelis et al., 2024; Ding et al., 2024; Yue et al., 2025; Chen et al., 2024c) or merging mechanisms to develop more versatile models (Yadav et al., 2024; Yu et al., 2024; Zhang et al., 2025b). Beyond aggregating static knowledge from multiple agents, multi-agent LLM training can also enhance reasoning capabilities. For example, multi-agent debates can generate diverse synthetic data, which can subsequently be used for supervised fine-tuning (Estornell et al., 2024; Li et al., 2024; Motwani et al., 2024; Dong & Ma, 2025; Perez et al., 2022; Ye et al., 2025a; Subramaniam et al., 2025). Reinforcement learning (RL) methods have also been adopted to improve LLM reasoning in areas such as alignment (Perez et al., 2022; Ma et al., 2023) and legibility (Kirchner et al., 2024a). Motwani et al. (2024) utilize a three-agent system for generation and fine-tune the models using Direct Preference Optimization (DPO). Reinforcement Learning with Generative Reward Models (GenRM) (Mahan et al., 2024; Ye et al., 2025b; Jiao et al., 2024; Wang et al., 2024b) represents another common approach of multi-agent training, where the reward signal is derived from the token probabilities of another LLM, coupled with the reasoning process. While our work aligns with these efforts, it diverges by using an additional tunable LLM to provide metacognitive instructions, guiding the low-level LLM during learning, rather than relying on a static GenRM. The most closely related works to ours are MAPoRL (Park et al., 2025) and COPRY (Ma et al., 2024). MAPoRL is a multi-agent debating framework that uses multi-agent reinforcement learning (MARL) with a learned verifier to fine-tune each LLM agent. COPRY duplicates an LLM into two agents, training them simultaneously in the roles of pioneer and observer using RL. However, unlike our approach, which explicitly separates metacognition from plan execution, these methods do not decompose the reasoning process but instead focus on improving direct chain-of-thought generation. Furthermore, our experiments are conducted on a larger scale and include more challenging problems.
5.3 Hierarchical Reasoning
Partitioning reasoning into hierarchical processes has been explored in prior research to make biological sense (Ye et al., 2018; Langley et al., 2004). In the context of language models, a hierarchical structure has been used to facilitate diverse reasoning patterns, including planning (Puerta-Merino et al., 2025; Sun et al., 2024; Song et al., 2023; Rana et al., 2023; Chen et al., 2024d; Yan et al., 2023; Xiao et al., 2024), validation (Haji et al., 2024; Xi et al., 2024) and self-refinement (Madaan et al., 2023; Kumar et al., 2024; Welleck et al., 2022). For instance, EvalPlanner (Saha et al., 2025b) is a framework that conducts reasoning through plan generation and execution. DOTS (Yue et al., 2024) extends decomposition by integrating a tree-based searching method with Analysis, Solution, and Verification layers. Marco-o1 (Zhao et al., 2024) focuses on open-ended problem-solving and abstract thinking, dynamically adjusting reasoning granularity and incorporating reflection mechanisms to enhance reasoning performance. Beyond these approaches, metacognition (Flavell, 1979) has been identified as another critical component of reasoning, referring to the intuitive understanding of one’s own cognitive and reasoning processes (Gao et al., 2024; Wang & Zhao, 2023). Wang & Zhao (2023) proposed a metacognitive prompting strategy to improve large language model (LLM) capabilities. Didolkar et al. (2024) further developed a prompt-guided method that enables models to label math problems with the required skills and subsequently use these labels to solve new problems. Gao et al. (2024) introduce meta-reasoner which use contextual multi-arm bandit to learn a high-level “advisor” over low-level reasoning process. Xiang et al. (2025) provides a Meta-CoT framework to think about its own thinking. They use construction-based methods as well as reinforcement learning to develop meta-cognitive skills. In our work, we also value reflecting on reasoning processes, and we enhance metacognitive abilities through two-agent interaction and reinforcement learning at both end.

### v3

@@ 3.2.1 Optimizing Single-turn MAMRP
where $R_{h}$ and $R_{l}$ are policies’ individual reward functions, including $R$ and regularization according to tasks and models, e.g., different format rewards (refer to Appendix C.2 for details). The detailed algorithm is in the Algorithm 1. We illustrate the MAMRP inference procedure and the proposed training method in Fig. 2. We also provide an analysis of different loss functions in Appendix C.5.
Figure 2: Comparison of training pipelines. Left: RL training of VRP and MRP, where a single LM agent is updated either with mixed (VRP) or explicit (MRP) meta-thinking. Middle: ReMA with separate parameters for the high-level (meta-thinking) and low-level (reasoning) agents; training alternates between freezing one agent and updating the other. Right: ReMA with shared parameters and multi-turn interactions: both agents share the same parameters and are distinguished by their system prompts. Training employs a turn-level ratio for stable multi-turn reinforcement learning and efficient updates, ensuring each turn’s contribution is controlled to prevent instability.
3.2.2 Scaling up to Multi-turn MAMRP
To scale up to multi-turn MAMRP, we can still adopt the iterative training strategy in Sec. 3.2.1. However, we make some changes to improve the efficiency of rollout and training.
First, we implement a parameter-sharing strategy where both high-level and low-level agents utilize identical model weights $\theta$, distinguished only by role-specific system prompts $S_{h}$ and $S_{l}$. Formally, we define $\pi_{h}=\pi_{\theta}(\cdot|S_{h},\cdot)$ and $\pi_{l}=\pi_{\theta}(\cdot|S_{l},\cdot)$, sharing the same underlying parameters rather than maintaining separate model instances. This approach eliminates the need for frequent model swapping on GPU during rollout, avoiding inefficient wait times, while enabling larger batch sizes during training to simultaneously optimize policies for both meta-thinking and reasoning roles.
Second, we propose a multi-turn GRPO with turn-level ratio to address the challenges of multi-turn MAMRP. The trajectory-level averaged objective with turn-level ratio of $\pi_{l}$ is defined as (The objective of $\pi_{h}$ is the similar but with different system prompt):
$\mathcal{J}(\theta)$ $=\mathbb{E}_{(\mathbf{x},\mathbf{y}^{*})\sim\mathcal{D},\,\{(\mathbf{m}_{i},\mathbf{y}_{i})\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|\mathbf{x})}$ (12)
$\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\frac{1}{|\mathbf{y}_{i,t}|}\sum_{j=1}^{|\mathbf{y}_{i,t}|}\Big{(}\min({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{i,t}(\theta)}\hat{A}_{i,t,j},\,\mathrm{clip}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{i,t}(\theta)},1-\epsilon,1+\epsilon)\hat{A}_{i,t,j})-\beta D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})\Big{)}\Bigg{]}$
where $\mathbf{y}_{i,t,j}$ is the $j$-th token at turn $t$ of the reasoning agent of the $i$-th trajectory. And the turn-level ratio for clipping is defined as:
$r_{i,t}(\theta)$ $=\frac{1}{|\mathbf{y}_{i,t}|}\sum\limits_{j=1}^{|\mathbf{y}_{i,t}|}r_{i,t,j}(\theta)=\frac{1}{|\mathbf{y}_{i,t}|}\sum\limits_{j=1}^{|\mathbf{y}_{i,t}|}\frac{\pi_{\theta}\bigl{(}\mathbf{y}_{i,t,j}\mid\mathbf{x},\{\mathbf{m}_{i,\cdot},\mathbf{y}_{i,\cdot}\}_{<t},\mathbf{m}_{i,t},\mathbf{y}_{i,t,<j}\bigr{)}}{\pi_{\theta_{\mathrm{old}}}\bigl{(}\mathbf{y}_{i,t,j}\mid\mathbf{x},\{\mathbf{m}_{i,\cdot},\mathbf{y}_{i,\cdot}\}_{<t},\mathbf{m}_{i,t},\mathbf{y}_{i,t,<j}\bigr{)}}.$ (13)
The introduction of the turn-level ratio serves two key purposes. First, using a token-level ratio (Eq. (2)) in the objective introduces bias for multi-turn training, as it averages over all tokens in a trajectory. This means that tokens within longer turns (those containing more tokens) can disproportionately influence the overall loss, and averaging at the token level may encourage excessively long single-turn responses. Second, clipping each token independently risks instability during training.
In contrast, the turn-level ratio aligns more closely with the underlying MDP formulation by treating all tokens within a turn as a single action and applying clipping at the turn level. Intuitively, this approach stabilizes training by preventing the LLM from making unstable updates that could result in extreme outputs, such as overly long repetitions or incoherent text. We conduct experimental verification in subsequent empirical results (Sec. 4.3).
@@ 4 Experiments
To evaluate the effectiveness and efficiency of ReMA, we conduct experiments on challenging benchmarks for two types of tasks: mathematical reasoning and LLM-as-a-Judge with three different LLMs. Then, we investigate the models’ performance in both single‑ & multi‑turn settings. Finally, we provide ablation studies and qualitative analyses of our method.
4.1 Experiment Settings
We first analyze the single-turn case of ReMA, i.e., $T=1$. The high-level agent generates a complete meta-thinking trace in one shot, and the low-level agent follows the instructions and outputs the final results. Single-turn ReMA reduces stochasticity and training cost while our experiments show that it still provides meaningful performance gains.
Benchmarks
We conduct experiments on two types of tasks: mathematical reasoning and LLM-as-a-Judge. For mathematical reasoning experiments, we train models on 7.5k training samples in MATH (Hendrycks et al., 2021) and use MATH500 (Lightman et al., 2023) as the in-distribution test dataset. Additionally, we test the optimized models on out-of-distribution datasets: GSM8K (Cobbe et al., 2021), AIME24111https://huggingface.co/datasets/AI-MO/aimo-validation-aime, AMC23222https://huggingface.co/datasets/AI-MO/aimo-validation-amc, GaoKao2023En (Zhang et al., 2023), Minerva Math (Lewkowycz et al., 2022), and Olympiad Bench (He et al., 2024).
For LLM-as-a-Judge benchmarks, we train models on RewardBench (Lambert et al., 2024). Specifically, we convert the original data into a pair-ranking format and split it into a training set of 5k items and a test set of 970 items, denoted as RewardBench970. The models are also tested on JudgeBench (Tan et al., 2024) to assess out-of-distribution performance. We refer to Appendix D.1.2 for detailed comparisons and results.
Baselines, Models, Training Settings
[…]
We train and test Llama-3-8B-Instruct, Llama-3.1-8B-Instruct (Dubey et al., 2024), and Qwen2.5-7B-Instruct (Team, 2024) on mathematical reasoning benchmarks. For LLM-as-a-judge benchmarks, we train and test Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct. We use instruct-tuned LLMs to prompt them to perform VRP, MRP, and MAMRP directly during training. Unless specified, we use two separate copies of the same model for high- and low-level agents in ReMA. We use the base reward setting in Appendix C.2 by default. And for the underlying RL algorithm, we use REINFORCE++ (Hu, 2025). We refer to Appendix D for detailed training settings.
4.2 Results of Single-turn ReMA
Question 1. Does single-turn ReMA outperforms baselines on both in-distribution and out-of-distribution test sets?
Table 1 compares the greedy decoding performance of ReMA against various RL baselines across mathematical benchmarks (Table 1(a)) and LLM-as-a-Judge benchmarks (Table 1(b)). Results across different LLMs indicate that, on average, ReMA outperforms all baselines, achieving a maximum improvement of 6.68% on mathematical benchmarks and 8.49% on LLM-as-a-Judge benchmarks.
Notably, ReMA achieves the highest performance on most benchmarks, particularly on out-of-distribution datasets, with a maximum improvement of 20% on AMC23 for Llama3-8B-Instruct, 13.33% on AIME24 for Qwen2.5-7B-Instruct, 14.23% on RewardBench970 for Llama3.1-8B-Instruct. These results demonstrate the superior out-of-distribution generalization ability conferred by the meta-thinking mechanism in ReMA. However, we observe that the accuracy gains from RL training on instruction-tuned LMs are smaller than from base models (Sec. 4.2.1). This may be due to the higher initial performance and the relatively fixed output distribution of instruction-tuned models, which limits the improvement and peak performance in RL.
@@ 4.2.1 Meta-thoughts boost low-level generalization
Figure 3: An RL experiment with 3 training schemes. While RL from SFT excels on easier problems, RL under Meta-thinking shows superior generalization to harder problems like AIME24.
Question 2. Can Reasoning benefit from Meta-thinking?
@@
Here we provide a tiny but motivating example of how ReMA gives better learning dynamics. We use Qwen2.5-Math-7B (Yang et al., 2024) as the starting base model, MATH (level 3-5, about 5.5K number of instances) as the training dataset, and we compare three reinforcement learning training schemes, in particular: (1) RL from Base: train the base model directly on MATH with binary outcome reward; (2) RL from SFT: SFT the base model with GPT-4o’s CoT answers; then RL on train dataset with binary outcome reward; (3) RL under Meta-thinking: SFT the base model with GPT-4o’s meta-thinking plans; then RL on train dataset with binary outcome reward.
The models are evaluated on 3 benchmarks (Fig. 3). SFT brings the best initial accuracy on in-distribution and easier sets, but fails to improve on harder ones. RL from Base yields limited gains. In contrast, RL under Meta-thinking achieves the best learning dynamics and generalizes better to challenging problems (AIME24). See Appendix F.1 for case studies.
4.2.2 Diverse meta-thinking characteristics of LLMs
Question 3. How well can LLMs learn to meta-think?
To further analyze meta-thinking behaviors, we train models with structured JSON-format actions inspired by Yue et al. . The meta-thinking agent generate two entries in one LM call, first selects from three actions: DECOMPOSE (breaking into subproblems), REWRITE (simplifying the problem), or EMPTY (direct solving), then generates the corresponding text. We compare Llama-3.1-8B-Instruct and Llama-3.2-1B-Instruct to study scale effects (two 1B models vs two 8B models) on meta-thinking agent’s training. We use vLLM guided JSON decoding (Dong et al., 2024) for valid formatting and base reward (reasoning agent’s solution accuracy with format constraints).
We observe that smaller LMs produce simpler outputs, likely due to limited capacity to maintain valid JSON formatting while exploring diverse reasoning strategies. As Fig. 4 shows, smaller LMs like Llama-3.2-1B-Instruct quickly converge to the simplest EMPTY action to avoid formatting penalties, while larger LMs like Llama-3.1-8B-Instruct can adapt meta-thinking strategies based on problem difficulty. See Appendix F.3 for detailed case studies.
Figure 4: Average problem difficulty by action type during training. Left: 1B LM collapses to the EMPTY action. Right: 8B LM adapts to a more complex meta-thinking strategy for harder problems.
4.3 Extending ReMA to Multi-turn MAMRP
We further extend ReMA to multi-turn MAMRP settings, enabling multiple rounds of interaction between the meta-thinking agent and the reasoning agent as defined in Eq. (7).
Unlike the inherent VRP capabilities of most LLMs, multi-turn ReMA requires initial bootstrapping. Thus, we constructed a supervised fine-tuning dataset (about 0.8K samples) from LIMO (Ye et al., 2025c) using GPT-4o to establish the starting point for multi-turn interaction capabilities. Then we finetune Qwen2.5-7B before RL training.
As described in Sec.3.2.2, we deploy the proposed GRPO with turn-level ratio clipping and trajectory-level averaging loss during training. And we remove the KL-divergence term to allow more flexible exploration. By default, the agents share the same parameters and are simultaneously updated using their trajectories. We refer to details in Appendix D.2.
Figure 5: Training results of multi-turn ReMA on MATH-Level3-5-8K under different rollout configurations.
Figure 6: Ablations of multi-turn ReMA on a tiny subset of MATH, we only show here the training curves of different training and rollout configurations.
4.3.1 Results and Ablations
Question 4. Can ReMA be scaled to multi-turn settings?
There are two key points revealed by our multi-turn ReMA experiments, as shown in Fig. 5. On one hand, the algorithm can demonstrate effective convergence on the training set, with accuracy steadily increasing from approximately 55% to 70% during training. It also achieves an average performance gain of about 5% across all seven test benchmarks, indicating stable improvements on out-of-distribution data. (Experiment with the rollout config of turn30_token512, see Appendix D.2.2 and Fig. 8 for more details.)
On the other hand, we observe that the performance of multi-turn ReMA is highly sensitive to hyperparameters such as the maximum response length per turn and the maximum number of turns. For certain configurations, the model either collapses into producing massive repetitions within a single turn or generates empty responses after only a few turns. Similar phenomena have been reported in concurrent works such as RAGEN (Wang et al., 2025), where these issues are attributed to the lack of fine-grained, reasoning-aware guidance. As a result, multi-turn RL becomes susceptible to long-horizon credit assignment challenges and state drift, often leading to reduced exploration diversity—a phenomenon referred to as the “Echo Trap”. To address this challenge, it is essential to comprehensively explore the training recipe w.r.t. model, data, and algorithm.
Question 5. How does parameter sharing and turn-level ratio affect multi-turn ReMA?
As shown in Fig. 6, we compare different configurations on a smaller dataset consisting of 133 samples—19 from each of the 7 MATH problem types—to evaluate sample efficiency and convergence speed. First, all configurations eventually achieve nearly 100% accuracy on the training dataset. Notably, the trajectory-level loss with turn-level ratio (Turn-Ratio, Eq. (13)) demonstrates substantially better sample efficiency than its token-level variants (Eq. (2)), reaching higher training rewards with fewer steps. We also present the training curve of separate weight setting, the empirical results show that shared parameters with simultaneous updates converge noticeably faster.